OpenStudy (anonymous):

Best Answer If you explain with answer! Which of the following is not equivalent to x - y + z? x + z - y z + x - y y - x + z z - y + x

OpenStudy (amoodarya):

take x=7 y=3 z=1 and check ! x - y + z=7-3+1=5 x + z - y =7+1-3=5 z + x - y=1+7-3=5 y - x + z=3-7+1=-3 z - y + x=1-3+7=5

OpenStudy (anonymous):

Thank you!